OpenStudy (anonymous):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves eight percent sales tax. Use a percent proportion to determine how much he would have to spend to save ten dollars. Which of the following is true? The solution is eighty dollars, and is a denominator in the proportion. The solution is eighty dollars, and is a numerator in the proportion. The solution is one hundred twenty-five dollars, and is a denominator in the proportion. The solution is one hundred twenty-five dollars, and is a numerator in the proportion.

OpenStudy (texaschic101):

Basically, you question is asking : 8% of what is 10 your percent proportion is : is/of = %/100 is = 10 % = 8 of = x now sub 10/x = 8/100 cross multiply 8x = 1000 x = 1000/8 x = 125 answer is D

OpenStudy (anonymous):

what d?

OpenStudy (texaschic101):

I meant the last answer....I just labeled your answer choices A,B,C, and D.